Prepare a schedule of budgeted cash collections from sales for May, June, and July. Include a three-month summary of estimated cash collections.Cash Receipts The sales budget for Andrew Inc. is forecasted as follows: Month Sales Revenue May $ 150,000 June 175,000 July
Answer #1

Solution :

May June July Total
Sales 150,000 175,000 160,000 485,000
60%os current month sale 90,000 105,000 96,000 291,000
Less: 2% discounts 1800 2100 1920 5,820
20% of previous month sale 30,000 30,000 35000 95,000
15% of second previous month sale 18,750 22,500 22500 63,750
136,950 155,400 151580 443,930

Working notes for the above answer

April sales collection remaining=40%
April sales=60000/40% 150,000
20% cash collection in May=150,000*20% 30,000
March Sales collection remaining =20%
March sales=25,000/20% 125,000
15% cash collection in =125,000*15% 18,750

Please do upvote if my answer helps you

Thank you !!